GlueX TOF Meeting, September 6, 2016

From GlueXWiki
Jump to: navigation, search

GlueX Time-of-Flight Meeting
Tuesday, September 6, 2016
11:00 am EDT
JLab: CEBAF Center, Room F326/327


  1. Announcements
  2. Review of minutes from the previous meeting
  3. TOF Analysis Status
    • Analysis of HV scan runs: 10838-10845 and 10646,10648
    • 4 hours of data taking HV scan provides sufficient data for gain calibration and find HV set points
    • In the current run plan 6 hours are foreseen.
  4. Run Plan
    • Which HV settings to be used (range, steps)
  5. Calibration Updates
  6. NIM paper



Talks can be deposited in the directory /group/halld/www/halldweb/html/talks/2016 on the JLab CUE. This directory is accessible from the web at .



  • JLab: Eugene Chudakov, Mark Ito (chair), David Lawrence, Simon Taylor, Beni Zihlmann

High Intensity Running Issues with the TOF

David showed slides illustrating the need to reduce the data volume from the TOF at high rate. The slides were from the Level 3 review.

The per crate limit is 1 Gbit/s as presently configured. We could upgrade the NIC to give 10 Gbit/s, but the VME backplane would limit us at about 1.5 Gbit/s. Crates TOF1 (with FADCs) and TOF2 (with TDCs) are two of the big contributors relative to other crates in the system. The amount of data can be reduced by tightening timing windows, enough to operate at high rates from a bandwidth perspective. For the TDCs the window is currently several microseconds wide. The reduction possible for the FADCs from time window tightening is not as great. Beni is afraid that raising the sparsification threshold may not do to much good; a lot of the hits are real particles. To get to high rate, we may have to upgrade the NIC for the flashes.

Eugene pointed out that the central part of the TOF may not be usable at high rate anyway. That needs to be studied, and alternative approaches explored.

Review of minutes from the previous meeting

We briefly looked at the minutes from August 23. Brad Cannon is meeting with Mark and Beni weekly now, with the exception of this week; he is away on travel.

TOF Analysis Status and Run Plan

Beni reviewed progress and plans for developing new high voltage set-points.

  • The analysis of HV scan runs, 10838-10845, 10646, and 10648 is now complete. The outer counters look to be efficient at 1200 V or perhaps a bit less. The picture from the inner counters is not as clear, but the low voltage points in the scan show voltage settings comparable to those of the outer counters.
  • The four hours of data used by the HV scan provided sufficient data for gain calibration and determining HV set points. We have asked for six hours for this run. We should run at nominal beam intensity to see if any problems with rate are still present at lower voltage.
  • Beni plans to run at a variety of fixed voltages for all tubes to simplify operations. Recall that with the Spring 2016 scan, voltages were adjusted with reference to their nominal value.

CAEN TDC calibration

Beni reminded us that we have not heard any recent news about the calibration to correct for the CAEN TDC system clock running at 41.7 MHz rather than 40 MHz. We need to look into what the plans are for this activity and schedule it.